# Channel Knowledge Map-assisted Dual-domain Tracking and Predictive Beamforming for High-Mobility Wireless Networks

### v1

@@ IV CKM-assisted C-Domain Tracking
In this section, we propose a CKM-assisted coordinated domain EKF tracking scheme that is applicable for both LoS-present and LoS-absent cases.
@@ IV-B C-Domain Measurement Model for LoS-present Case
If the LoS path exists in the $[\eta]$-th slot, the C-Domain measurements consist of the AoA, delay, and Doppler shift of the LoS path, which is denoted with the subscript ${}^{\prime\prime}1^{\prime\prime}$. Based on the geometric relationships depicted in Fig. 1, the C-Domain measurement model is given by
$\begin{cases}&\tau_{1}[\eta]=\frac{2||\boldsymbol{q}[\eta]||_{2}}{c}+w_{\tau}[\eta]\in\mathbb{R},\\ &\mu_{1}[\eta]=-\frac{2v_{1}[\eta]q_{y}[\eta]f_{c}}{c||\boldsymbol{q}[\eta]||_{2}}+w_{\mu}[\eta]\in\mathbb{R},\\ &\cos(\theta_{1}[\eta])=\frac{q_{x}[\eta]}{||\boldsymbol{q}[\eta]||_{2}}+w_{\cos\theta}[\eta]\in(-1,1),\end{cases}$ (4)
@@ IV-D EKF-based C-Domain Tracking
Due to the non-linear measurement model for both the LoS-present case in (5) and LoS-absent case in (6), we adopt the EKF [34] to track the state of the vehicle in the C-Domain. In the $\eta$-slot, assuming we have the knowledge of the state vector ${\boldsymbol{\alpha}}[\eta]\in\mathbb{R}^{3\times 1}$ estimated at the $\eta$-slot, the state prediction for the $[\eta+1]$-slot is given by
@@
and the prediction covariance matrix $\boldsymbol{C}_{r}[\eta+1|\eta]\in\mathbb{R}^{3\times 3}$, $r=1,2$ is given by $\boldsymbol{C}_{r}[\eta+1|\eta]=\boldsymbol{E}\boldsymbol{C}_{r}[\eta]\boldsymbol{E}^{T}+\boldsymbol{Q}_{\alpha}$, where $\boldsymbol{C}_{r}[\eta]$ is the posterior covariance matrix for the C-Domain state update in the $[\eta]$-th slot.
@@ IV-D1 LoS-present case
To update the state in the $[\eta]$-th time slot, the EKF employs a linearized measurement model around the predicted state [^6] In our notation, only the state prediction and state update are denoted with a hat (e.g., $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ and $\hat{\boldsymbol{\alpha}}[\eta]$); other variables, such as the measurement $\boldsymbol{\beta}[\eta|\eta-1]$ and $\boldsymbol{\beta}[\eta]$, are presented without a hat.. Specifically, we have: $\boldsymbol{\beta}_{1}[\eta]\approx g_{1}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{1}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{1}}[\eta]$, where $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ is the predicted C-Domain state in $[\eta-1]$-th time slot, $\boldsymbol{G}_{1}\in\mathbb{R}^{3\times 3}$ represents the Jacobian matrix of $g_{1}$ with respect to the state vector and it is given by (8)
@@
$=\left.\begin{bmatrix}\frac{2q_{x}[\eta]}{c||\boldsymbol{q}[\eta]||_{2}}&\frac{2q_{y}[\eta]}{c||\boldsymbol{q}[\eta]||_{2}}&0\\ -\frac{2v_{1}[\eta]f_{c}q_{y}[\eta]q_{x}[\eta]}{c||\boldsymbol{q}[\eta]||_{2}^{3}}&\frac{2f_{c}v_{1}[\eta](q_{x}[\eta])^{2}}{c||\boldsymbol{q}[\eta]||_{2}^{3}}&-\frac{2q_{y}[\eta]f_{c}}{c||\boldsymbol{q}[\eta]||_{2}}\\ \frac{q_{y}[\eta]^{2}}{||\boldsymbol{q}[\eta]||_{2}^{3}}&-\frac{q_{x}[\eta]q_{y}[\eta]}{||\boldsymbol{q}[\eta]||_{2}^{3}}&0\end{bmatrix}\right|_{\hat{\boldsymbol{\alpha}}[\eta|\eta-1]}.$
@@ IV-D2 LoS-absent case
where $g_{2}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])\in\mathbb{R}^{3P\times 1}$ can be obtained via referring to CKM based on the predicted vehicle state. The Jacobian matrix $\boldsymbol{G}_{2}$ represents the partial derivatives of the measurement function $g_{2}(\cdot)$ with respect to the C-Domain state vector and is computed numerically as follow:
@@
Following the standard EKF framework [18], the procedures for proposed C-Domain tracking in both LoS-present $(r=1)$ and LoS-absent $(r=2)$ scenarios are outlined below: (1) Linearization: $\boldsymbol{\beta}_{r}[\eta]\approx g_{r}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{r}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{r}}[\eta]$, (2) Kalman gain matrix: $\boldsymbol{K}_{r}=\boldsymbol{C}_{r}[\eta|\eta-1]\boldsymbol{G}_{r}^{T}(\boldsymbol{G}_{r}\boldsymbol{C}_{r}[\eta|\eta-1]\boldsymbol{G}_{r}^{T}+\boldsymbol{Q}_{\beta_{r}})^{-1}$, (3) State update: $\hat{\boldsymbol{\alpha}}[\eta]=\hat{\boldsymbol{\alpha}}[\eta|\eta-1]+\boldsymbol{K}_{r}(\boldsymbol{\beta}_{r}[\eta]-g_{r}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1]))$, (4) Posterior covariance matrix: $\boldsymbol{C}_{r}[\eta]=(\boldsymbol{I}_{3}-\boldsymbol{K}_{r}\boldsymbol{G}_{r})\boldsymbol{C}_{r}[\eta|\eta-1]$, (5) State prediction: $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]=\boldsymbol{E}\hat{\boldsymbol{\alpha}}[\eta]$, (6) Prediction covariance matrix: $\boldsymbol{C}_{r}[\eta+1|\eta]=\boldsymbol{E}\boldsymbol{C}_{r}[\eta]\boldsymbol{E}^{T}+\boldsymbol{Q}_{\alpha}$. Interested readers can refer to [18, 9, 31, 3] for further details of EKF. The proposed EKF scheme requires identifying whether the LoS is present or not, a process already proposed in [4] and [9]. The computational complexity of the proposed EKF scheme is constant, i.e., $\mathcal{O}(1)$, in the LoS-present case, whereas in the LoS-absent case, it increases to $\mathcal{O}(P^{3})$ due to the inversion of a $3P\times 3P$ matrix required in the Kalman gain computation.
@@ V-A B-Domain State Evolution Model
Similar to [35] and [6], define $\Theta\triangleq\{\tilde{\theta}_{\kappa}={\pi(\kappa-1)}/{N_{\theta}}\}^{N_{\theta}}_{\kappa=1}$ as an angular grid with ${N}_{\theta}$ equal segments spanning the entire angular space $[0,\pi)$. Also, the B-Domain state evolution model is defined as a discrete Markov process, i.e.,
$\small\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])=\boldsymbol{p}(\theta_{i}[\eta])\boldsymbol{\Pi}_{i}[\eta]\in[0,1]^{1\times N_{\theta}},$ (11)
@@
where $\boldsymbol{\Pi}_{i}[\eta]\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ denotes the TPM for the AoA of $i$-th path in the $[\eta]$-th time slot, $\boldsymbol{p}({\theta}_{i}[\eta])$ and $\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])$ denote the probability mass function (PMF) of the AoA for the $i$-th path in the $[\eta]$-th slot and the predicted PMF of the AoA of $i$-th path in the $[\eta+1]$-th time slot, respectively. The $(\kappa,\iota)$-th element of the TPM for the $i$-th path in the $[\eta]$-th slot is given by $\boldsymbol{\Pi}_{i}[\eta]_{\kappa,\iota}=\operatorname{Pr}\{\theta_{i}[\eta+1]=\tilde{\theta}_{\iota}\mid\theta_{i}[\eta]=\tilde{\theta}_{\kappa}\},\forall\kappa,\iota\in\{1,\dots,N_{\theta}\}$, where $\operatorname{Pr}\{\theta_{i}[\eta+1]=\tilde{\theta}_{\iota}\mid\theta_{i}[\eta]=\tilde{\theta}_{\kappa}\}$ denotes the transition probability of the AoA of the $i$-th path from angle $\tilde{\theta}_{\kappa}$ at time slot $[\eta]$ to $\tilde{\theta}_{\iota}$ at time slot $[\eta+1]$. Clearly, the TPM should satisfy the following requirements: (1) ${0}\leq\boldsymbol{\Pi}_{i}[\eta]_{\kappa,\iota}\leq{1}$ for $\forall\kappa,\iota\in\{1,\dots,N_{\theta}\}$. (2) The sum of each row of $\boldsymbol{\Pi}_{i}[\eta]$ should be equal to 1, i.e., $||\boldsymbol{\Pi}_{i}[\eta]_{\kappa,:}||_{1}=1,\forall\kappa\in\{1,\dots,N_{\theta}\}$.
@@
Both the channel temporal correlation and the CKM provide a prior information on beam transition and thus affect the TPM. Therefore, the angular TPM for the $i$-th path at $[\eta]$-th slot is modeled as:
$\small\boldsymbol{\Pi}_{i}[\eta]=(1-c_{\Pi}[\eta])\cdot\boldsymbol{\Pi}_{i}^{I}[\eta]+c_{\Pi}[\eta]\cdot\boldsymbol{\Pi}_{i}^{II}[\eta],$ (12)
@@
where $\boldsymbol{\Pi}_{i}^{I}[\eta]$ denotes the TPM caused by the channel temporal correlation and $\boldsymbol{\Pi}_{i}^{II}[\eta]$ denotes the TPM designed using the channel a priori information provided by CKM. Real constant $c_{\Pi}[\eta]\in[0,1]$ is a weight parameter that is tuned to characterize the contribution of a prior information provided by channel temporal correlation and CKM. The value of $c_{\Pi}[\eta]$ depends on the condition of beam transition, the received echo SNR, and the horizontal speed of the target vehicle relative to the RSU.
@@
As mentioned earlier, the AoA of the $i$-th path between two consecutive time slots should remain close even in high-mobility wireless networks, provided there is no random birth/death of paths. Following [35] and [36], the $(\kappa,\iota)$-th element of $\boldsymbol{\Pi}_{i}^{I}$ of a path in the $[\eta]$-th slot is given by
@@
$\small\left\{\boldsymbol{\Pi}_{i}^{I}[\eta]\right\}_{\kappa,\iota}=\begin{cases}\zeta\xi^{\left|\kappa-\iota\right|},&{\text{if}}\left|\kappa-\iota\right|\leq\epsilon,\\ {0},&{\text{otherwise,}}\end{cases}$ (13)
@@
where $\forall i\in\{1,\dots,P\}$. The parameter $\xi\in[0,1]$ characterizes the variation speed of angles over time, where $\xi\rightarrow 0$ indicates almost no change and $\xi\rightarrow 1$ indicates significant change. The variable $\epsilon$ represents the boundary of AoA variation and $\zeta$ is a normalization coefficient as in [17, 6]. Both $\xi$ and $\epsilon$ depend on the relative speed of the vehicle to the RSU, given by $\sin(\hat{\theta}_{1}[\eta+1|\eta])\hat{v}_{1}[\eta+1|\eta]$. On the other hand, the $(\kappa,\iota)$-th element of $\boldsymbol{\Pi}_{i}^{II}$ in the $[\eta]$-th slot is given by
@@
$\small\left\{\boldsymbol{\Pi}_{i}^{II}[\eta]\right\}_{\kappa,\iota}=\varsigma\exp\left(-\frac{(\hat{\theta}_{i}[\eta+1|\eta]-\tilde{\theta_{\iota}})^{2}}{2\sigma_{\text{CKM}}^{2}}\right),\forall\kappa,$ (14)
@@
  • Stationary Transition $(I)$: In this case, both $\boldsymbol{\Pi}_{i}^{I}$ and $\boldsymbol{\Pi}_{i}^{II}$ are available due to the time-correlated variations in AoA and the CKM mapping between the position of vehicle and AoA. In this case, $c_{\Pi}\in(0,1)$.
  • Nonstationary but Predictable Transition $(II)$: The angle may abruptly change due to the fixed scatterers in the environment which have been recorded in the CKM (shown by the yellow arrows in the figure). In this case, the angle transition is not time-correlated and $\boldsymbol{\Pi}_{i}^{I}$ can not be exploited. Fortunately, CKM has stored this beam transition and thus the angle transition is still predictable based on $\boldsymbol{\Pi}_{i}^{II}$. In this case $c_{\Pi}$ is equal to 1.
  • Nonstationary and Unpredictable Transition $(III)$: The random block is caused by randomly appearing scatterers/obstacles in the environment (illustrated by the red asterisks in the figure). For this case, AoA prediction is impossible and no a priori information can be used. In this case, a confirm distribution can be adopted, i.e., $\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])=\frac{1}{N_{\theta}}\cdot\boldsymbol{1}_{N_{\theta}}$.

### v2

@@ III CKM-assisted Dual-domain Tracking Framework
Remark 1: It is worth noting that the proposed dual-domain framework can be naturally extended to the joint tracking of other channel parameters, such as delay and Doppler shifts. Specifically, the B-Domain state vector can be extended to include the delay and Doppler shift of each path in addition to AoA. The temporal state evolution can then be modeled by constructing TPMs analogous to those developed for the AoA parameter, and a higher-dimensional CKM that encodes the joint distribution of multiple channel parameters needs to be established. However, this extension comes at the cost of increased computational complexity in the B-Domain tracking.
@@ IV CKM-assisted C-Domain Tracking
In this section, we propose a CKM-assisted coordinated domain EKF tracking scheme that is applicable for both LoS-present and LoS-absent cases. The proposed C-Domain tracking scheme requires identifying whether the LoS is present or not, a problem already studied in [53] and [45].
@@ IV-B C-Domain Measurement Model for LoS-present Case
If the LoS path exists in the $\eta]-$th slot, the C-Domain measurements consist of the AoA, delay, and Doppler shift of the LoS path, which is denoted with the subscript $^{\prime\prime}1^{\prime\prime}$. Based on the geometric relationships depicted in Fig. 1, the C-Domain measurement model is given by
[…]
@@ IV-D EKF-based C-Domain Tracking
Due to the non-linear measurement model for both the LoS-present case in (5) and LoS-absent case in (6), we adopt the EKF [16] to track the state of the vehicle in the C-Domain. In the $\eta$-slot, assuming we have the knowledge of the state vector ${\boldsymbol{\alpha}}[\eta]\in\mathbb{R}^{3\times 1}$ estimated at the $\eta$-slot, the state prediction for the $\eta+1]-$slot is given by
@@
and the prediction covariance matrix $\boldsymbol{C}_{r}[\eta+1|\eta]\in\mathbb{R}^{3\times 3}$, $r=1,2$ is given by $\boldsymbol{C}_{r}[\eta+1|\eta]=\boldsymbol{E}\boldsymbol{C}_{r}[\eta]\boldsymbol{E}^{T}+\boldsymbol{Q}_{\alpha}$, where $\boldsymbol{C}_{r}[\eta]$ is the posterior covariance matrix for the C-Domain state update in the $\eta]-$th slot.
@@ IV-D1 LoS-present case
To update the state in the $\eta]-$th time slot, the EKF employs a linearized measurement model around the predicted state [^6] In our notation, only the state prediction and state update are denoted with a hat (e.g., $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ and $\hat{\boldsymbol{\alpha}}[\eta]$); other variables, such as the measurement $\boldsymbol{\beta}[\eta|\eta-1]$ and $\boldsymbol{\beta}[\eta]$, are presented without a hat.. Specifically, we have: $\boldsymbol{\beta}_{1}[\eta]\approx g_{1}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{1}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{1}}[\eta]$, where $\hat{\boldsymbol{\alpha}}[\eta|\eta-1]$ is the predicted C-Domain state in $\eta-1]-$th time slot, $\boldsymbol{G}_{1}\in\mathbb{R}^{3\times 3}$ represents the Jacobian matrix of $g_{1}$ with respect to the state vector and it is given by (8)
@@
[…]
@@ IV-D2 LoS-absent case
where $g_{2}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])\in\mathbb{R}^{3P\times 1}$ can be obtained via referring to CKM based on the predicted vehicle state. In the LoS-absent case, the AoA of NLoS paths can vary significantly with slight vehicle movement due to multi-path propagation. Consequently, a high spatial resolution CKM is essential for capturing such angular variations. The Jacobian matrix $\boldsymbol{G}_{2}$ represents the partial derivatives of the measurement function $g_{2}(\cdot)$ with respect to the C-Domain state vector and is computed numerically as follow:
@@
Following the standard EKF framework [24], the procedures for proposed C-Domain tracking in both LoS-present $(r=1)$ and LoS-absent $(r=2)$ scenarios are outlined below: (1) Linearization: $\boldsymbol{\beta}_{r}[\eta]\approx g_{r}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{G}_{r}(\boldsymbol{\alpha}[\eta]-\hat{\boldsymbol{\alpha}}[\eta|\eta-1])+\boldsymbol{w}_{\beta_{r}}[\eta]$, (2) Kalman gain matrix: $\boldsymbol{K}_{r}=\boldsymbol{C}_{r}[\eta|\eta-1]\boldsymbol{G}_{r}^{T}(\boldsymbol{G}_{r}\boldsymbol{C}_{r}[\eta|\eta-1]\boldsymbol{G}_{r}^{T}+\boldsymbol{Q}_{\beta_{r}})^{-1}$, (3) State update: $\hat{\boldsymbol{\alpha}}[\eta]=\hat{\boldsymbol{\alpha}}[\eta|\eta-1]+\boldsymbol{K}_{r}(\boldsymbol{\beta}_{r}[\eta]-g_{r}(\hat{\boldsymbol{\alpha}}[\eta|\eta-1]))$, (4) Posterior covariance matrix: $\boldsymbol{C}_{r}[\eta]=(\boldsymbol{I}_{3}-\boldsymbol{K}_{r}\boldsymbol{G}_{r})\boldsymbol{C}_{r}[\eta|\eta-1]$, (5) State prediction: $\hat{\boldsymbol{\alpha}}[\eta+1|\eta]=\boldsymbol{E}\hat{\boldsymbol{\alpha}}[\eta]$, (6) Prediction covariance matrix: $\boldsymbol{C}_{r}[\eta+1|\eta]=\boldsymbol{E}\boldsymbol{C}_{r}[\eta]\boldsymbol{E}^{T}+\boldsymbol{Q}_{\alpha}$. Interested readers can refer to [24, 45, 38, 36] for further details of EKF. Note that $\boldsymbol{\beta}_{r}[\eta]$ is always obtained from the received signals. In the LoS-absent case, the measurement model is characterized by the CKM, which provides a numerical mapping from the target state to the multipath parameters, thereby ensuring continuous and stable state update. The computational complexity of the proposed EKF scheme is constant, i.e., $\mathcal{O}(1)$, in the LoS-present case, whereas in the LoS-absent case, it increases to $\mathcal{O}(P^{3})$ due to the inversion of a $3P\times 3P$ matrix required in the Kalman gain computation.
@@ V-A B-Domain State Evolution Model
Similar to [51] and [50], define $\Theta\triangleq\{\tilde{\theta}_{\kappa}={\pi(\kappa-1)}/{N_{\theta}}\}^{N_{\theta}}_{\kappa=1}$ as an angular grid with ${N}_{\theta}$ equal segments spanning the entire angular space $0,\pi)$. Also, the B-Domain state evolution model is defined as a discrete Markov process, i.e.,
$\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])=\boldsymbol{p}(\theta_{i}[\eta])\boldsymbol{\Pi}_{i}[\eta]\in[0,1]^{1\times N_{\theta}},$ (11)
@@
where $\boldsymbol{\Pi}_{i}[\eta]\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ denotes the TPM for the AoA of $i$-th path in the $\eta]-$th time slot, $\boldsymbol{p}({\theta}_{i}[\eta])$ and $\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])$ denote the probability mass function (PMF) of the AoA for the $i$-th path in the $\eta]-$th slot and the predicted PMF of the AoA of $i$-th path in the $\eta+1]-$th time slot, respectively. The $(\kappa,\iota)$-th element of the TPM for the $i$-th path in the $\eta]-$th slot is given by $\boldsymbol{\Pi}_{i}[\eta]_{\kappa,\iota}=\operatorname{Pr}\{\theta_{i}[\eta+1]=\tilde{\theta}_{\iota}\mid\theta_{i}[\eta]=\tilde{\theta}_{\kappa}\},\forall\kappa,\iota\in\{1,\dots,N_{\theta}\}$, where $\operatorname{Pr}\{\theta_{i}[\eta+1]=\tilde{\theta}_{\iota}\mid\theta_{i}[\eta]=\tilde{\theta}_{\kappa}\}$ denotes the transition probability of the AoA of the $i$-th path from angle $\tilde{\theta}_{\kappa}$ at time slot $\eta]$ to $\tilde{\theta}_{\iota}$ at time slot $\eta+1]$. Clearly, the TPM should satisfy the following requirements: (1) ${0}\leq\boldsymbol{\Pi}_{i}[\eta]_{\kappa,\iota}\leq{1}$ for $\forall\kappa,\iota\in\{1,\dots,N_{\theta}\}$. (2) The sum of each row of $\boldsymbol{\Pi}_{i}[\eta]$ should be equal to 1, i.e., $\boldsymbol{\Pi}_{i}[\eta]_{\kappa,:}||_{1}=1,\forall\kappa\in\{1,\dots,N_{\theta}\}$.
@@
Both the channel temporal correlation and the CKM provide a prior information on beam transition and thus affect the TPM. We model the fusion of these two sources as a convex combination of TPMs and such a convex combination remains mathematically valid, as established in [27, 1]. Therefore, the angular TPM for the $i$-th path at $\eta]-$th slot is modeled as:
$\boldsymbol{\Pi}_{i}[\eta]=(1-c_{\Pi_{i}}[\eta])\cdot\boldsymbol{\Pi}_{i}^{I}[\eta]+c_{\Pi_{i}}[\eta]\cdot\boldsymbol{\Pi}_{i}^{II}[\eta],$ (12)
@@
where $\boldsymbol{\Pi}_{i}^{I}[\eta]$ denotes the TPM caused by the channel temporal correlation and $\boldsymbol{\Pi}_{i}^{II}[\eta]$ denotes the TPM designed using the channel a priori information provided by CKM. The real constant $c_{\Pi_{i}}[\eta]\in[0,1]$ is the probability of selecting the transition mechanism $\boldsymbol{\Pi}_{i}^{II}[\eta]$ provided by CKM, while $(1-c_{\Pi_{i}}[\eta])$ corresponds to the probability of selecting the transition mechanism $\boldsymbol{\Pi}_{i}^{I}[\eta]$ provided by channel temporal correlation. The value of $c_{\Pi_{i}}[\eta]$ depends on the condition of beam transition, the received echo SNR, the horizontal speed of the target vehicle relative to the RSU, and the quality of the constructed CKM.
@@
As mentioned earlier, the AoA of the $i$-th path between two consecutive time slots should remain close even in high-mobility wireless networks, provided there is no random birth/death of paths. Following [51] and [2], the $(\kappa,\iota)$-th element of $\boldsymbol{\Pi}_{i}^{I}$ of a path in the $\eta]-$th slot is given by
@@
$\left\{\boldsymbol{\Pi}_{i}^{I}[\eta]\right\}_{\kappa,\iota}=\begin{cases}\zeta\xi^{\left|\kappa-\iota\right|},&{\text{if}}\left|\kappa-\iota\right|\leq\epsilon,\\ {0},&{\text{otherwise,}}\end{cases}$ (13)
@@
where $\forall i\in\{1,\dots,P\}$. The parameter $\xi\in[0,1]$ characterizes the variation speed of angles over time, where $\xi\rightarrow 0$ indicates almost no change and $\xi\rightarrow 1$ indicates significant change. The variable $\epsilon$ represents the boundary of AoA variation and $\zeta$ is a normalization coefficient as in [30, 50]. Both $\xi$ and $\epsilon$ depend on the relative speed of the vehicle to the RSU, given by $\sin(\hat{\theta}_{1}[\eta+1|\eta])\hat{v}_{1}[\eta+1|\eta]$. On the other hand, the $(\kappa,\iota)$-th element of $\boldsymbol{\Pi}_{i}^{II}$ in the $\eta]-$th slot is given by
@@
$\left\{\boldsymbol{\Pi}_{i}^{II}[\eta]\right\}_{\kappa,\iota}=\varsigma\exp\left(-\frac{(\hat{\theta}_{i}[\eta+1|\eta]-\tilde{\theta_{\iota}})^{2}}{2\sigma_{\text{CKM}}^{2}}\right),\forall\kappa,$ (14)
@@
  • Stationary Transition $(I)$: In this case, both $\boldsymbol{\Pi}_{i}^{I}$ and $\boldsymbol{\Pi}_{i}^{II}$ are available due to the time-correlated variations in AoA and the CKM mapping between the position of vehicle and AoA. In this case, $c_{\Pi_{i}}\in(0,1)$. A larger $c_{\Pi_{i}}$ is preferred under a low received echo SNR or the target vehicle traveling with a high horizontal speed relative to the RSU, where the weaker channel temporal correlation or unreliable measurements necessitate greater reliance on the CKM prior. In contrast, a lower quality constructed CKM shifts reliance toward temporal channel correlation, leading to a smaller $c_{\Pi_{i}}$.
  • Non-stationary but Predictable Transition $(II)$: The angle may abruptly change due to the fixed scatterers in the environment which have been recorded in the CKM (shown by the yellow arrows in the figure). In this case, the angle transition is not time-correlated and $\boldsymbol{\Pi}_{i}^{I}$ can not be exploited. Fortunately, CKM has stored this beam transition and thus the angle transition is still predictable based on $\boldsymbol{\Pi}_{i}^{II}$. In this case $c_{\Pi_{i}}$ is equal to 1.
  • Non-stationary and Unpredictable Transition $(III)$: The random block is caused by randomly appearing scatterers/obstacles in the environment (illustrated by the red asterisks in the figure). For this case, AoA prediction is impossible and no a priori information can be used. In this case, a confirm distribution can be adopted, i.e., $\boldsymbol{p}({\theta}_{i}[\eta+1|\eta])=\frac{1}{N_{\theta}}\cdot\boldsymbol{1}_{N_{\theta}}$.